OpenStudy (anonymous):

Simplify. 2/20 + 3/5 - 4/5

OpenStudy (anonymous):

(2/20)+(12/20)-(16/20) (making the denominator same) =>-(2/20) =>-(1/10)

OpenStudy (anonymous):

\[\frac{2}{20}+\frac{3}{5}-\frac{4}{5} = \frac{2}{20}+\frac{3*4}{5*4}-\frac{4*4}{5*4} = \frac{2}{20}+\frac{12}{20}-\frac{16}{20} = \frac{2+12-16}{20}=\frac{14-16}{20}\] =\[\frac{-2}{20}=\frac{-1}{10}\]

OpenStudy (anonymous):

thank you both :)

OpenStudy (anonymous):

2/20+3/5-4/5 2/20+3#4/5*4-4*4/5*4 -2/20 -1/10

OpenStudy (anonymous):

i mean - not #